OpenStudy (anonymous):

 The sum of two consecutive odd integers is 36. Find the integers

OpenStudy (anonymous):

x +x +2 =36 2x=34 x=17 x+2=19

OpenStudy (anonymous):

\[n+n+2=36\] \[2n+2=36\] \[2n=34\] \[n=17,n+2=19\]